Please help solve this
2e^x = 6
I need the value of x

Answer by yannick(6)


2e^x = 6
by dividing by 2 the both sides
(2e^x)/2 = 6/2 , we find e^x = 3
By using the definition e^log a = a,
we come out with the solution
x = log 3